TERRI'S WEEKLY INCOME IS A $100 DOLLARS LESS THAN DOULBE GARY'S WEEKLY IONCOME. IF TERRI MAKE'S $448 A WEEK, WHAT DOES GARY MAKE?

AND I HAVE A NOTHER QUESTION
PAUL MAKES $25 A WEEK LEES THAN THE SUM OF WHAT AVERY AND CASEY MAKES TOGETHER MAKE. CASEY'S WEEKLY INCOME WOULD BE TRIPLE STEVE'S IF CASEY MAKES $50 MORE A WEEK. PAUL MAKES $342 A WEEK. HOW MUCH D OES AVERY MAKE?

Answer by sdmmadam@yahoo.com(530)

TERRI'S WEEKLY INCOME IS A $100 DOLLARS LESS THAN DOULBE GARY'S WEEKLY IONCOME. IF TERRI MAKE'S $448 A WEEK, WHAT DOES GARY MAKE?
Let Gary make G dollars a week.
Given that Terry's weekly income is $448
Given that TERRI'S WEEKLY INCOME IS A $100 DOLLARS
LESS THAN DOULBE GARY'S WEEKLY IONCOME
That is 448 is $100 < 2G
That is 448 = 2G - 100
That is (448+100) = 2G
548 = 2G
Therefore G = 548/2 = 274
Therefore Gary's weekly income is 274 dollars
Verification: Terry' should be 100 less than twice Gary's
And 448 is of course 100 less than (2X274)= 548